# GigaTok: Scaling Visual Tokenizers to 3 Billion Parameters for Autoregressive Image Generation

### v1

@@ Abstract
00footnotetext: Work partly done as an Intern at ByteDance. Correspondence Author.
@@
Figure 2: The 2.9B GigaTok achieves SOTA autoregressive image generation with a 1.4B AR model on ImageNet 256$\times$256 resolution.
@@ 2 Related Work
Image tokenizers. Image tokenizers map image inputs into discrete or continuous tokens which can be modeled by downstream generative models. Continuous tokenizers are built upon Variational Autoencoders (VAE) [28, 29], and discrete tokenizers such as VQ-GAN [15] quantize visual features into discrete visual tokens during training. Vector Quantization (VQ) [15, 56, 65] is dominantly adopted for discretizing tokens, while other quantization methods [68, 74, 49, 75] focus on scaling codebook size for better tokenizers. However, how to properly scale up tokenizer models has rarely been studied in existing literature. A concurrent work ViTok [75] attempts to scale continuous VAE-based tokenizers for downstream diffusion models, but ends up suggesting de-prioritizing tokenizer scaling due to its less predictable effect for downstream models. In contrast to previous work, we provide a detailed analysis of the reconstruction vs. generation dilemma for scaling tokenizers and the solution to it.
Autoregressive Visual Generation. Autoregressive visual generation models [15, 65, 33, 50, 40, 49, 56, 38, 59, 58] follow the next-token-prediction approach of LLMs, enabling them to leverage advancements in LLMs and simplifying the path to unified multi-modal generation. Other methods using discrete tokenizers incorporate visual-specific paradigms, such as mask image modeling [8, 60, 69, 68] and next-scale-prediction [53, 36], for improved performance. We focus on the autoregressive paradigm and reveal that scaling tokenizers helps AR models to be comparable to the best of those visual generation models.
Semantic Guidance for Visual Models. Recent work has explored the use of guidance from visual foundation models [7, 43, 46, 71, 23] to enhance compressed visual latents with richer semantics [36, 9, 75, 18, 62, 67, 72, 64, 63, 10, 76], as well as to improve the representations in visual generation models [70]. In contrast to these approaches, we analyze the dilemma in scaling tokenizers and emphasize the critical role of semantic regularization in their effective scaling.
@@ 3.1 AR Probing for Tokenizer Evaluation
Specifically, we use the tokenizer to train a small Llama-style model [54, 50] ( [^1] M parameters) for 50 epochs, and evaluate its gFID [24], validation loss, and linear probing accuracy [11, 23] for a fair comparison between different tokenizers. Training the proposed AR Probing model for evaluating tokenizers is 10$\times$ more efficient than training the original 343M downstream AR model. Our experiments in Sec. 5.1 (Fig. 6) demonstrate that the trends observed with AR Probing align with the performance of the large-scale AR models after sufficient training.
@@
Linear probing accuracy. Beyond visual generation quality, we also investigate whether scaling tokenizers will lead to better visual representations learned by the AR models, which may provide inspiration for future research in unified multimodal understanding and generation with AR models. To assess the representation quality of AR models, we adopt the standard practice [11, 23] of evaluating linear probing accuracy using features from the middle Transformer layer of the AR probing model.
@@ 3.2 Naively Scaling Tokenizers Does Not Work
To study the challenges when naively scaling visual tokenizers, we train three vector-quantized tokenizers [^1] The tokenizer architectures are described in Sec. 4.1 on ImageNet [48] at 256$\times$256 resolution with increasing model sizes. As shown in Fig. 3, as the tokenizer size increases, although the reconstruction quality (rFID) consistently improves, the AR generation performance (gFID) significantly degrades. This highlights the reconstruction vs. generation dilemma in tokenizer scaling. Moreover, we observe that the validation loss of AR Probing consistently increases as the tokenizers scale, indicating that larger tokenizers lead to complicated token dependencies that are more difficult for the AR model to learn. This observation motivates us to design the semantic regularization to constrain the latent space complexity of the tokenizer and therefore break the reconstruction vs. generation dilemma in Sec. 4.2.
@@ 4.1 Architecture
In current literature, the CNN [32] architectures are the dominant choices for image tokenizers [15, 40, 68, 75] due to their effectiveness in capturing fine-grained local details. Yet, Transformers are more scalable architectures with less inductive bias. Thus, we design a vector quantized (VQ) tokenizer backbone with a hybrid architecture that combines CNN [32, 15] and Transformer [57, 13, 6] for encoder and decoder (Fig. 4). Specifically, our encoder consists of a series of CNN blocks which progressively downsamples the input image by a factor of $p$, followed by Transformer layers and a vector quantizer to produce discrete latent codes. Similarly, our decoder consists of multiple Transformer layers, followed by CNN decoders which upsamples the features to obtain the reconstructed image [^2] Throughout this work, we use downsample ratio $p=16$, codebook dimension $D=8$, and codebook size $16384$ by default.. Our tokenizer architecture can be adapted to both 1D and 2D tokenizers by utilizing different Transformer designs introduced in the next two paragraphs.
@@ 4.3 Entropy Loss for Billion-Level Tokenizers
$\mathcal{L}_{\text{entropy}}=\mathbb{E}_{\mathbf{z}}\left[H(\hat{\mathbf{z}}|\mathbf{z})\right]-H(\hat{\mathbf{z}})$ (3)
where $H(\cdot)$ denotes the Shannon entropy, $\mathbf{z}\in\mathbb{R}^{D}$ is the input for quantizer to be quantized to $\mathbf{\hat{z}}=\mathbf{c}_{i}\in\mathbb{R}^{D}$ and $\mathbf{c}_{i}$ is the $i$-th codebook vector. $\mathbb{E}_{\mathbf{z}}\left[H(\hat{\mathbf{z}}|\mathbf{z})\right]$ penalizes the uncertainty in quantization to reduce quantization error, and $-H(\hat{\mathbf{z}})$ encourages the codebook vectors to be selected more uniformly across the entire codebook. The detailed derivation can be found in our supp. We find that the entropy penalty addresses the convergence difficulty of large tokenizers. As shown in Fig. 5, introducing entropy loss to the 2.9B tokenizer enables the codebook usage to quickly reach a high level, and the loss converges properly [^3] We take perceptual loss as an example, and reconstruction loss shows a similar pattern.
@@ 5.2 Scaling with Semantic Regularization
Table 3: System-level comparison for tokenizers and downstream generation models on ImageNet 256$\times$256. For gFID, we present the lowest value between w/ or w/o CFG scenarios. †: Training set includes data besides ImageNet. ‡: Using frozen DINO [7] for discriminator, which largely improves rFID. $\star$: Without classifier-free-guidance. $\diamond$: Data from BiGR [22].
@@ 5.5 Discussion and Ablation Study
Table 5: Layer $l$ for semantic regularization (S-S tokenizer). Smaller $l$ brings better downstream AR model representations but can sacrifice reconstruction and downstream generation quality. We choose $l$=3 by default for more balanced performance.
@@
Discussion on generation costs. When generating an image, AR models take multiple passes to predict tokens, while tokenizers only need one forward pass. Therefore, the time consumption for decoding tokens to images is relatively small compared to AR models. We record the ratio of time spent on tokenizer decoding for different tokenizer/AR models in Tab. 4. For a 1.4B AR model, our largest 2.9B tokenizer takes only $\sim$10of the total inference time.
@@
Exploring pretrained semantic encoder choices. We compare CLIP-B (DFN) [16, 46], SigLIP-400M [71] and DINOv2-B [43] as the source of semantic regularization for S-B tokenizers. As shown in Tab. 6, utilizing DINOv2-B as the semantic encoder for regularization produces the best tokenizer for reconstruction, downstream class conditional generation and representation quality.
@@ 6 Conclusion
7 Acknowledgments
The authors sincerely thank Qihang Yu and Liang-Chieh Chen for their valuable discussions during the development of GigaTok.
@@ Appendix C Detailed GigaTok Implementation
where $l_{2}(\mathbf{z},\mathbf{C})\in\mathbb{R}^{N}$ is the $L_{2}$ distance between $\mathbf{z}$ and all the codebook vectors. Then, minimization of the quantization error can be partially achieved by minimizing the expectation of entropy $\mathbb{E}_{\mathbf{z}}\left[H(\hat{\mathbf{z}}|\mathbf{z})\right]$, which can be understood as maximizing the prediction confidence for $p(\hat{\mathbf{z}}|\mathbf{z})$. To encourage higher codebook utilization, we aim to make the average appearance probability of codebook vectors more uniform. This is achieved by maximizing the entropy $H(\hat{\mathbf{z}})$, Therefore, the optimization of the two entropy terms leads to the final entropy loss equation:
$\mathcal{L}_{\text{entropy}}=\mathbb{E}_{\mathbf{z}}\left[H(\hat{\mathbf{z}}|\mathbf{z})\right]-H(\hat{\mathbf{z}})$ (5)
In practice, to calculate $H(\hat{\mathbf{z}})$, we estimate $p(\hat{\mathbf{z}}=\mathbf{c}_{i})$ by $p(\hat{\mathbf{z}}=\mathbf{c}_{i})=\mathbb{E}_{\mathbf{z}}\left[p(\hat{\mathbf{z}}=\mathbf{c}_{i}|\mathbf{z})\right]$. Note that entropy loss is not our contribution. We only provide a detailed definition of entropy loss in VQ scenarios for better understanding.
@@
Table 10: Full results for our tokenizers and AR models on ImageNet 256$\times$256. For gFID, we present the lowest value between w/ or w/o CFG scenarios. ‡: Using frozen DINO [7] for discriminator, which largely improves rFID. $\star$: Without classifier-free-guidance.
@@ Appendix G More Discussions About Related Work
ViTok [21] is a concurrent work which has explored the effect of model scaling for VAE [28]. ViTok evaluates its VAE models in terms of both reconstruction and downstream diffusion generation performance. While having a very different setting from our GigaTok, ViTok similarly finds that asymmetric design is better for VAEs. While ViTok suggests that small encoders are optimal, we point out that in our setting scaling encoders is also beneficial. Notably, the reconstruction vs. generation dilemma for scaling visual tokenizers is not solved in ViTok. We hypothesize that adding semantic regularization may similarly help solve the tokenizer scaling dilemma for VAEs, but leave it for future study.
MAGVIT-v2 [68] introduces LFQ to enhance discrete tokenizers. It also introduces the entropy penalty for tokenizer training, which is shown to be important for training large-scale tokenizers in our work. Instead of tokenizer model scaling, MAGVIT-v2 focuses more on scaling the codebook size of tokenizers. While codebook dimension and codebook size are important bottlenecks for visual tokenizers, we point out that model size scaling is also an important aspect.
ImageFolder [36] utilizes two branches for image encoding to handle high-level semantic information and low-level visual details respectively. It reveals the importance of the learned representation of tokenizers.

### v2

@@ Abstract
00footnotetext: Corresponding Author.
@@
Figure 2: The 2.9B GigaTok achieves SOTA autoregressive image generation with a 1.4B AR model on ImageNet $\times256$ resolution.
@@ 2 Related Work
Image tokenizers. Image tokenizers map image inputs into discrete [15, 56, 66] or continuous [28] tokens which can be modeled by downstream generative models. For discrete tokenizers, Vector Quantization (VQ) [15, 56, 66] is dominantly adopted. Recently, new quantization methods [69, 75, 49, 76] have also been proposed for better scaling of codebook size. However, how to properly scale up tokenizer models is insufficiently studied in existing literature. ViT-VQGAN [66] and TiTok [70] utilize transformer architecture to enable convenient scaling of tokenizers, but end up training their best generative models on smaller tokenizer versions. A concurrent work, ViTok [76], suggests de-prioritizing VAE scaling due to its less predictable effect for downstream diffusion models. We observe a similar reconstruction vs. generation dilemma in scaling discrete tokenizers, and provide our analysis and solution to it.
Autoregressive Visual Generation. Autoregressive visual generative models [66, 33, 50, 40, 49, 56, 38, 60, 58] follow the next-token-prediction (NTP) approach of LLMs, enabling the leverage of advancements in LLMs and simplifying the path to unified multi-modal generation. Other variants utilize visual-specific paradigms such as mask image modeling [8, 61, 70, 69] and next-scale-prediction [53, 36] for better performance. We reveal that scaling tokenizers helps NTP AR models to be comparable to these variants.
Semantic Guidance for Visual Generative Models and Tokenizers. The guidance from visual foundation models [7, 43, 46, 72, 23] has been used to improve training convergence speed and quality [71, 65] of visual generative models, as well as enhancing representation quality or downstream performance of visual tokenizers [36, 9, 76, 18, 63, 68, 73, 65, 64, 10, 77, 59, 41]. REPA [71] presents impressive performance improvements brought by a simple representation alignment strategy, and recently, VA-VAE [65] shows the significant benefits of semantic guidance to the reconstruction-generation Pareto Frontier of VAEs. Different from existing work, GigaTok novelly reveals the critical role of semantic regularization for resolving the reconstruction vs. generation dilemma in scaling visual tokenizers.
@@ 3.1 AR Probing for Tokenizer Evaluation
[…]
@@
Linear probing accuracy. Beyond visual generation quality, we also investigate whether scaling tokenizers will lead to better visual representations of AR models, which may provide inspiration for future research in unified multimodal understanding and generation with AR models. To assess the representation quality, we adopt the standard practice [11, 23] of linear probing accuracy using features from the middle Transformer layer of the AR probing model.
@@ 3.2 Naively Scaling Tokenizers Does Not Work
To study the challenges when naively scaling visual tokenizers, we train three vector-quantized tokenizers [^1] The tokenizer architectures are described in Sec. 4.1 on ImageNet [48] at $\times256$ resolution with increasing model sizes. As shown in Fig. 3, as the tokenizer size increases, although the reconstruction quality (rFID) consistently improves, the AR generation performance (gFID) significantly degrades. This highlights the reconstruction vs. generation dilemma in tokenizer scaling. Moreover, we observe that the validation loss of AR Probing consistently increases as the tokenizers scale, indicating that larger tokenizers lead to complicated token dependencies that are more difficult for the AR model to learn. This observation motivates us to design the semantic regularization to constrain the latent space complexity of the tokenizer and therefore break the reconstruction vs. generation dilemma in Sec. 4.2.
@@ 4.1 Architecture
The CNN [32] architectures have been the dominant choices for image tokenizers [15, 40, 69, 76] due to their effectiveness in capturing fine-grained local details. Yet, Transformers are more scalable architectures with less inductive bias. Thus, we design a vector quantized tokenizer backbone with a hybrid architecture that combines CNN [32, 15] and Transformer [57, 13, 6] for encoder and decoder (Fig. 4). Specifically, our encoder consists of a series of CNN blocks that progressively downsamples the input image by a factor of $p$, followed by Transformer layers and a vector quantizer to produce discrete latent codes. Similarly, our decoder consists of multiple Transformer layers, followed by CNN decoders which upsamples the features to obtain the reconstructed image [^2] Throughout this work, we use downsample ratio $p=16$, codebook dimension $D=8$, and codebook size $16384$ by default.. Our tokenizer architecture can be adapted to both 1D and 2D tokenizers by using different Transformer designs introduced in the next two paragraphs.
@@ 4.3 Entropy Loss for Billion-Level Tokenizers
[…]
@@ 5.2 Scaling with Semantic Regularization
Table 3: System-level comparison for tokenizers and downstream generation models on ImageNet $\times256$. For gFID, we present the lowest value between w/ or w/o CFG scenarios. †: Training set includes data besides ImageNet. ‡: Using frozen DINO [7] for discriminator, which largely improves rFID. $\star$: Without classifier-free-guidance. $\diamond$: Data from BiGR [22].
@@ 5.5 Discussion and Ablation Study
Table 5: Layer $l$ for semantic regularization (S-S tokenizer). Smaller $l$ brings better downstream AR model representations but can sacrifice reconstruction and downstream generation quality. We choose ll=3 by default for more balanced performance.
@@
Discussion on generation costs. When generating an image, AR models take multiple passes to predict tokens, while tokenizers only need one forward pass. Therefore, the time consumption for decoding tokens to images is relatively small compared to AR models. We record the ratio of time spent on tokenizer decoding for different tokenizer/AR models in Tab. 4. For a 1.4B AR model, our largest 2.9B tokenizer takes only $\sim$10% of the total inference time.
@@
Exploring pretrained semantic encoder choices. We compare CLIP (DFN) [16, 46], SigLIP-400M [72] and DINOv2-B [43] as the source of semantic regularization for S-B tokenizers. As shown in Tab. 6, utilizing DINOv2-B as the semantic encoder for regularization produces the best tokenizer for reconstruction, downstream class conditional generation and representation quality.
@@ 6 Conclusion
Acknowledgments
This work is partially supported by the National Nature Science Foundation of China (No. 62402406).
The authors also sincerely thank Qihang Yu and Liang-Chieh Chen for their valuable discussions during the development of GigaTok.
@@ Appendix C Detailed GigaTok Implementation
[…]
@@
Table 10: Full results for our tokenizers and AR models on ImageNet $\times256$. For gFID, we present the lowest value between w/ or w/o CFG scenarios. ‡: Using frozen DINO [7] for discriminator, which largely improves rFID. $\star$: Without classifier-free-guidance.
@@ Appendix G More Discussions About Related Work
ViTok [21] is a concurrent work which has explored the effect of model scaling for VAE [28]. ViTok evaluates its VAE models in terms of both reconstruction and downstream diffusion generation performance. While having a very different setting from GigaTok, ViTok similarly finds that asymmetric design is better for VAEs. While ViTok suggests that small encoders are optimal, we point out that in our setting scaling encoders is also beneficial. Notably, the reconstruction vs. generation dilemma for scaling visual tokenizers is not solved in ViTok. We hypothesize that adding semantic regularization may similarly help solve the tokenizer scaling dilemma for VAEs, but leave it for future study.
MAGVIT-v2 [69] introduces LFQ to enhance discrete tokenizers. It also introduces the entropy penalty for tokenizer training, which is shown to be important for training large-scale tokenizers in our work. Instead of tokenizer model scaling, MAGVIT-v2 focuses more on scaling the codebook size of tokenizers. While codebook dimension and codebook size are important bottlenecks for visual tokenizers, we point out that model size scaling is also an important way for improving visual tokenizers.
ImageFolder [36] utilizes two branches for image encoding to handle high-level semantic information and low-level visual details respectively. It seminally utilizes semantic alignment to enhance the learned representation of tokenizers.
VA-VAE [65] tames the reconstruction vs. generation dilemma in increasing latent dimensions for continuous VAE [28, 29]. VA-VAE improves the reconstruction-generation Pareto Frontier by introducing vision foundation model alignment loss. In contrast, we seek continuous improvements in both reconstruction and generation by scaling tokenizers. Semantic regularization serves different purposes in the two works.